OpenStudy (anonymous):

What is the vertical asymptote of 3x-1/2x+9

Directrix (directrix):

Is this supposed to be an equation? The equation, y = (3x - 1) / (2x + 9) ? @Fifi11

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Don't I just do 2x+9 = 0 and solve for it?

Directrix (directrix):

Solve for x in 2x+9 = 0, yes. That will give you the value of x which cause the rational function to be undefined. Note: Check to make sure that 2x + 9 and 3x -1 have no common factors.

Directrix (directrix):

@Fifi11 Let me know what you get as an answer and I will tell you what I got. Okay?

OpenStudy (anonymous):

-9/2s

Directrix (directrix):

>>Don't I just do 2x+9 = 0 and solve for it? 2x + 9 = 0 2x + 9 - 9 = 0 - 9 Solve this for x. --> @Fifi11 I did not see an "s" in the equation.

OpenStudy (anonymous):

Yeah it's -9/2 I meant to say negative nine halfs

Directrix (directrix):

The vertical asymptote is an equation. So, to call it -9/2 is incorrect. What is the equation of the vertical asymptote?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."